OpenStudy (karlaltr):

A student records the internal temperature of a hot sandwich that has been left to cool on a kitchen counter, The room temperature is 90 degrees celcius. An equation that models this situation is T (t) = 63(0.5)^t/10 +19 How much time did it take for the sandwich to reach an internal temperature of 30 degrees celcius?

OpenStudy (karlaltr):

30 = 63 (0.5) ^ t/10 + 19 then what?

OpenStudy (itsmichelle29):

You have to go backwards :)

OpenStudy (itsmichelle29):

Add 10+19 and so on

OpenStudy (karlaltr):

but its not 19 + 10 its a fraction of t/10

OpenStudy (itsmichelle29):

CAN U DRAW THE EQUATION OUT?

OpenStudy (karlaltr):



OpenStudy (dumbcow):

start by isolating the exponential part \[0.5^t = \frac{30-19}{63}\] Then take natural log of both sides use log property where exponent can be brought to front \[t \ln(0.5) = \ln(\frac{11}{63})\] Finally solve for t \[t = \frac{\ln(\frac{11}{63})}{\ln(0.5)}\]